QUESTION 9 Find the Laplace Transform of f(t)= - 1 if ts 4; f(t) = 1 if t> 0.
Answer #1

There might be some mistake in the question. As it gives 2 values in the interval 0<t<4

The function might be given

f(t)=\left\{\begin{matrix} -1& \text{if } \ t\leq 4\\ 1 & \text{if } \ t>4 \end{matrix}\right.

\mathcal{L}(f)=\int_{0}^{\infty}f(t)\cdot e^{-st}dt

=\int_{0}^{4}f(t)\cdot e^{-st}dt+\int_{4}^{\infty}f(t)\cdot e^{-st}dt

=\int_{0}^{4}(-1)\cdot e^{-st}dt+\int_{4}^{\infty}1\cdot e^{-st}dt

=-\left [ -\frac{1}{s}e^{-st} \right ]_0^4+\left [ -\frac{1}{s}e^{-st} \right ]_4^{\infty}

=\frac{1}{s}e^{-4s}-\frac{1}{s}+\frac{1}{s}e^{-4s}

=\frac{2}{s}e^{-4s}-\frac{1}{s}

This is the required answer.